# Jaike van Twiller1    Yossiri Adulyasak2    Erick Delage2    Djordje Grbic1    Rune Møller Jensen1

### v3

@@ E.3 Hyperparameters
E.4 Additional Experiments
In Table 5, we analyze different configurations of feasibility regularization (FR). First, we evaluate performance under the same hyperparameters as AM-P policies. Second, we examine the effect of significantly increasing $\lambda_{f}$. Third, we assess performance with specific hyperparameter tuning, including $\lambda_{f}$. These experiments indicate that FR can reduce the distance to the feasible region, however, achieving fully feasible instances remains a challenge.
@@ E.4 Additional Experiments
Alg. Model F.M. H.P. $\lambda_{f}$ Ob. (\$) Time (s) F.I. (%) $d(\textit{PH}(s_{t}))$
SAC AM FR Proj. 0.0677 1139.78 13.82 0.00 62.77
SAC AM FR Proj. 0.677 1031.55 13.45 0.00 120.74
SAC AM FR Tune 0.065 1113.03 12.63 0.00 37.55
PPO AM FR Proj. 0.0677 2606.87 13.36 0.00 3171.86
PPO AM FR Proj. 0.677 2593.29 13.20 0.00 3381.59
PPO AM FR Tune 0.302 1842.46 11.74 0.00 754.21
Table 5: Performance evaluation on $N$ instances of feasibility regularization (FR) with hyperparameter (H.P.) settings: projected hyperparameters (Proj.), ensuring a fair comparison with projection-based policies, and tuned hyperparameter (Tune), optimized specifically for PPO and SAC with FR. While we use $\lambda_{f}$ as the control parameter for FR, tuning involves adjusting a Lagrangian multiplier for each constraint. Average performance metrics include objective value in profit (Ob.), inference time in seconds (Time), percentage of feasible instances (F.I.), and total absolute distance to the feasible region $d(\textit{PH}(s_{t}))$. Note that † indicates infeasible objectives.

### v5

@@ Appendix E Deep RL Implementation Details
Appendix F Additional Experiments
In Table 5, we analyze the vanilla AM policy with feasibility regularization (FR) with static Lagrangian multipliers ($\lambda_{f}$), and primal-dual methods (PD) that learn the Lagrangian multipliers ($\lambda_{f}$). We show the performance learned through SAC and PPO, and also show how post-hoc feasibility recovery (FR+R, PD+R) impacts the objective value.
@@ Appendix F Additional Experiments
Alg. Model F.M. Obj. (\$) Feas. (%) $d(\textit{PH}(s_{t}))$
SAC AM FR 1,188.90 0.00 83.03
SAC AM FR+R 1,217.22 100.00 0.00
SAC AM PD 1,077.84 0.00 30.21
SAC AM PD+R 1,080.21 100.00 0.00
PPO AM FR 1,294.31 0.00 2,565.51
PPO AM FR+R 1,133.18 100.00 0.00
PPO AM PD 1,149.44 0.00 1,441.97
PPO AM PD+R 885.85 100.00 0.00
Table 5: Performance on NN instances of feasibility regularization (FR), primal-dual methods (PD), and their feasibility-recovered variants (FR+R, PD+R). Metrics include average objective profit (Obj.), inference time (Time), feasibility rate (Feas.), and total absolute distance to the feasible region $d(\textit{PH}(s_{t}))$.